# On regions of mixed unitarity for semigroups of unital quantum channels

### v1

@@ Abstract.
It is established that both discrete and continuous semigroups of unital quantum channels are eventually mixed unitary. However, after introducing the mixed unitary index of a unital quantum channel as the least time beyond which all subsequent powers of the channel are mixed unitary, we demonstrate that for any fixed finite dimension $d\geq 3$, there exists no universal upper bound for this index. Furthermore, for a continuous semigroup that is not mixed unitary for some $t>0$, we prove it remains non-mixed unitary for all times $t>0$, sufficiently close to the origin. Finally, a necessary and sufficient condition is derived for a quantum dynamical semigroup to be a convex combination of maps implemented solely by Weyl unitaries.
@@
Completely positive maps, quantum channels, semigroup, mixed unitary channels.
@@ 1. Introduction
For a natural number $d$, let $\mathbb{M}_{d}$ denote the algebra of all $d\times d$ complex matrices. In this article, we will be restricting ourselves to studying unital (doubly-stochastic) quantum channels from $\mathbb{M}_{d}$ to $\mathbb{M}_{d}$. A unitary quantum channel on $\mathbb{M}_{d}$ is a quantum channel which is given by $X\mapsto U^{*}XU$, for all $X\in\mathbb{M}_{d}$ and for some unitary $U\in\mathbb{M}_{d}$. Unitary quantum channels are extreme in the convex set of unital quantum channels on $\mathbb{M}_{d}$. A quantum channel is said to be a mixed unitary if it is in the convex hull of the set of all unitary quantum channels, or equivalently (due to finite dimensionality) it is a convex combination of unitary channels. That is, a quantum channel $\Phi:\mathbb{M}_{d}\to\mathbb{M}_{d}$ is a mixed unitary if there exist $\{\lambda_{j}\}_{j=1}^{\ell}\subseteq[0,1]$ and unitaries $\{U_{j}\}_{j=1}^{\ell}\subseteq\mathbb{M}_{d}$ for some $l\in\mathbb{N}$ such that $\sum_{j=1}^{\ell}\lambda_{j}=1$ and
@@
Mixed unitary channels play a significant role in quantum information theory for several key reasons. They provide a rich class of examples that capture key properties of general quantum operations [Ros08]. By virtue of their structural simplicity, they remain highly useful for theoretical analysis, often proving more tractable than arbitrary quantum channels. These channels emerge naturally in multiple contexts, including majorization frameworks based on Hermitian operators [AU82], as well as in quantum cryptographic protocols involving state encryption[AMTB00, HDPW04, AS04].
Taking inspiration from the classical Birkhoff-von Neumann theorem, which realizes all doubly stochastic maps as convex combinations of permutations, one may suspect that perhaps every unital quantum channel is mixed unitary. A well-known result [LaSt93, Theorem 1] establishes this for $d=2$. However, this property fails for $d\geq 3$ as first demonstrated by Tregub [Tre86] and Kümmerer-Maassen [KuMa87]. See also [LaSt93, BPS93]. Additionally, determining whether a given unital channel is mixed unitary or not is known to be NP-hard [LWat20].
@@
$\delta_{d}(X):=\frac{\mbox{trace}(X)}{d}I_{d},\penalty 10000\ \penalty 10000\ X\in\mathbb{M}_{d},$
@@
A famous theorem of Watrous [Wat09], says that there is a ball of positive radius around $\delta_{d}$ such that all unital channels in it are mixed unitary. As a consequence, any sequence of unital channels converging to the completely depolarizing channel eventually consists of only mixed unitary channels. This was taken note of in the context of powers of primitive unital channels in Section 4 of [HRS20]. They define mixed unitary time (or index) of a unital quantum channel as a threshold $k$ such that all subsequent powers beyond which are mixed unitary. Moreover they show that such a threshold exists for primitive unital channels and obtain an upper bound for it.
@@
In a more recent work, Kribs et al., [KJRM23] extend this idea very significantly. Note that the completely depolarizing channel can be viewed as the unique conditional expectation map from $\mathbb{M}_{d}$ to the $C^{*}$-algebra $\mathbb{C}I_{d}$ of scalars. Fix a unital $C^{*}$-subalgebra $\mathcal{A}$ of $\mathbb{M}_{d}$ and let $\mathbf{E}_{\mathcal{A}}$ be the unique trace-preserving conditional expectation to it. It is a unital quantum channel. In [KJRM23] it is observed that unital quantum channels sufficiently close to $\mathbf{E}_{\mathcal{A}}$ and fixing elements of $\mathcal{A}$ are mixed unitary. Armed with this result they show that if $\Phi$ is a unital quantum channel on $\mathbb{M}_{d}$, then there exists some $k\in\mathbb{N}$ such that $\Phi^{k}$ is mixed unitary [KJRM23, Theorem 22]. However, $\Phi^{k}$ fixes elements of an algebra $\mathcal{A}$, does not imply that $\Phi^{k+1}$ does the same. For this reason, they couldn’t show that $\{\Phi^{n}\}_{n\geq 0}$ has finite mixed unitary time. We overcome this difficulty in Theorem 3.9, by making use of the notion of conditional automorphism on the peripheral eigenspace of $\Phi$.
This raises the natural question as to for fixed $d$, whether there exists a bound $N_{d}$ such that mixed unitary indices of all unital quantum channels on $\mathbb{M}_{d}$ are bounded by $N_{d}$. We answer this question in the negative in Theorem 4.9. We get it as a consequence of our analysis of continuous ( one parameter) semigroups of unital quantum channels.
@@
A deep study of continuous semigroups of mixed unitary channels was carried out long time ago by Kümmerer and Maassen in [KuMa87]. Their goal was to understand Markov dilations of such semigroups. In the process they characterized generators of quantum dynamical semigroups which remain mixed unitary for all times. We obtain a different characterization of such generators (Theorem 5.2). We show that even in continuous time semigroups of unital channels eventually become mixed unitary. In other words, they all have finite mixed unitary index (Theorem 4.7).
@@
We organize the paper as follows: In Section 2, we recall the necessary definitions and basic results required in the subsequent sections, and fix some notations. Section 3, investigates the mixed unitarity for a discrete semigroup $\{\Phi^{n}\}_{n\in\mathbb{N}}$ generated by an unital quantum channel $\Phi$ on $\mathbb{M}_{d}$. We first establish that every trace-preserving *-automorphism on a $C^{*}$-subalgebra $\mathcal{A}$ of $\mathbb{M}_{d}$ is necessarily of the form $\mathrm{Ad}_{U}$ for some $U\in\mathbb{M}_{d}$ (see Theorem 3.6). As a consequence of this Theorem, we prove that any unital quantum channel on $\mathbb{M}_{d}$ is asymptotically a conditional *-automorphism (Theorem 3.8). This leads us to one of the main results, namely every discrete semigroup $\{\Phi^{n}\}_{n\in\mathbb{N}}$ of unital quantum channels on $\mathbb{M}_{d}$ is eventually mixed unitary (Theorem 3.9).
@@
In section 4, we extend the analysis to continuous one parameter semigroups $\{\Phi_{t}\}_{t\geq 0}$ of unital quantum channels on $\mathbb{M}_{d}$. An analogous theorem is established, proving these semigroups are also eventually mixed unitary (Theorem 4.7). We also prove that there is no universal upper bound of the mixed unitary index for a fixed dimension (Theorem 4.9). That is, there is no $N\geq 1$ such that $\Phi^{N}$ is a mixed unitary channel, for all unital channels $\Phi$ on $\mathbb{M}_{d}$.
@@
In Section 5, we examine the local behavior of a continuous semigroup of a unital quantum channels $\{\Phi_{t}\}_{t\geq 0}$ near $t=0$. First, a characterization theorem is established for such a semigroup on $\mathbb{M}_{d}$ to be mixed unitary for all times (Theorem 5.2). Using this Theorem, we prove that every continuous one parameter semigroup $\{\Phi_{t}\}_{t\geq 0}$ of unital quantum channels on $\mathbb{M}_{d}$ is either mixed unitary for all times or there is a time $t_{0}$ such that $\Phi_{t}$ is never mixed unitary for all times $t\in(0,t_{0})$ (Theorem 5.4). In the last section, we discuss mixed Weyl-unitary channels on $\mathbb{M}_{d}$. We provide a complete structure theorem (Theorem 6.4) for a one-parameter semigroup $\{\Phi_{t}\}_{t\geq 0}$ of unital quantum channels on $\mathbb{M}_{d}$ to be mixed Weyl-unitary for all times. We also introduce the notion of $G$-mixed unitary channel on $\mathbb{M}_{d}$ for any subgroup $G$ of $\mathbb{U}(d)$. For any closed subgroup $G$, we have an analogous theorem (see Theorem 6.8) for a continuous one parameter semigroup $\{\Phi_{t}\}_{t\geq 0}$ of unital quantum channels on $\mathbb{M}_{d}$ to be $G$-mixed unitary for all times.
@@ 2.1. Notations
Let $\mathbb{C}^{d}$ denote the $d$-dimensional complex Hilbert space with the inner product linear in the second variable and conjugate linear in the first. The set $\{e_{i}:1\leq i\leq d\}$ is the standard orthonormal basis of $\mathbb{C}^{d}$. Let $\mathbb{M}_{d_{1}\times d_{2}}$ be the algebra of complex matrices of size $d_{1}\times d_{2}$, we use $\mathbb{M}_{d}$ to refer $\mathbb{M}_{d\times d}$. The identity matrix in $\mathbb{M}_{d}$ is denoted as $I=I_{d}$. Additionally, we let $\operatorname{\textsf{T}}=\operatorname{\textsf{T}}_{d}:\mathbb{M}_{d}\to\mathbb{M}_{d}$ be the transpose map (w.r.t. the standard basis), $\operatorname{\textsf{tr}}=\operatorname{\textsf{tr}}_{d}:\mathbb{M}_{d}\to\mathbb{C}$ be the trace map and $\operatorname{\textsf{id}}=\operatorname{\textsf{id}}_{d}:\mathbb{M}_{d}\to\mathbb{M}_{d}$ be the identity map. By writing $A=[a_{ij}]\in\mathbb{M}_{d_{1}\times d_{2}}$, we mean that $A$ is a complex matrix of size $d_{1}\times d_{2}$, with $(i,j)^{th}$ entry equal to $a_{ij}$, for all indices satisfying $1\leq i\leq d_{1}$ and $1\leq j\leq d_{2}$. If $A=[a_{ij}]\in\mathbb{M}_{d_{1}\times d_{2}}$, the adjoint of $A$, denoted as $A^{*}$, is a matrix $A^{*}=[b_{ij}]\in\mathbb{M}_{d_{2}\times d_{1}}$, where each entry $b_{ij}$ is given by $\overline{a_{ji}}$ for all $1\leq i\leq d_{2}$ and $1\leq j\leq d_{1}$. A matrix $A\in\mathbb{M}_{d}$ is said to be Hermitian if $A=A^{*}$, and a Hermitian matrix $A$ is said to be positive (semi-definite) if $\langle x,Ax\rangle\geq 0$ for all $x\in\mathbb{C}^{d}$. The collection of all positive matrices in $\mathbb{M}_{d}$ is denoted by $\mathbb{M}_{d}^{+}$. If $A\in\mathbb{M}_{d}^{+}$, then we write $A\geq 0$. A matrix $U\in\mathbb{M}_{d}$ is said to be unitary if $U^{*}U=UU^{*}=I$, and denotes the set of all $d\times d$ unitary matrices by $\mathbb{U}(d)$.
For every $x\in\mathbb{C}^{d_{1}}$ and $y\in\mathbb{C}^{d_{2}}$, define the rank one linear map $|{x}\rangle\langle{y}|:\mathbb{C}^{d_{2}}\to\mathbb{C}^{d_{1}}$ by $|{x}\rangle\langle{y}|(z):=\langle y,z\rangle x$, for all $z\in\mathbb{C}^{d_{2}}$. It is straightforward to verify that $A|{x}\rangle\langle{y}|B=|{Ax}\rangle\langle{B^{*}y}|$ holds for any $x\in\mathbb{C}^{d_{1}}$, $y\in\mathbb{C}^{d_{2}}$, and matrices $A\in\mathbb{M}_{d_{1}},B\in\mathbb{M}_{d_{2}}$. Additionally, it is easy to see that the matrix $|{x}\rangle\langle{x}|$ is positive for all $x\in\mathbb{C}^{d}$.
Given the linear maps $\Phi_{1}:\mathbb{M}_{d_{1}}\to\mathbb{M}_{d_{2}}$ and $\Phi_{2}:\mathbb{M}_{d_{3}}\to\mathbb{M}_{d_{4}}$, the tensor product map $\Phi_{1}\otimes\Phi_{2}:\mathbb{M}_{d_{1}}\otimes\mathbb{M}_{d_{3}}\to\mathbb{M}_{d_{2}}\otimes\mathbb{M}_{d_{4}}$ is the unique map determined by $\Phi_{1}\otimes\Phi_{2}(A\otimes B)=\Phi_{1}(A)\otimes\Phi_{2}(B)$ for all elements $A\in\mathbb{M}_{d_{1}}$ and $B\in\mathbb{M}_{d_{3}}$.
Definition 2.1.
A linear map $\Phi:\mathbb{M}_{d_{1}}\to\mathbb{M}_{d_{2}}$ is said to be
  1. (i)
    Hermitian preserving if $\Phi(X^{*})=\Phi(X)^{*}$ for all $X\in\mathbb{M}_{d_{1}}$;
  2. (ii)
    positive if $\Phi(\mathbb{M}_{d_{1}}^{+})\subseteq\mathbb{M}_{d_{2}}^{+}$;
  3. (iii)
    completely positive (CP) if the map $\operatorname{\textsf{id}}_{k}\otimes\Phi:\mathbb{M}_{k}\otimes\mathbb{M}_{d_{1}}\to\mathbb{M}_{k}\otimes\mathbb{M}_{d_{2}}$ is positive, for all $k\geq 1$;
  4. (iv)
    unital if $\Phi(I_{d_{1}})=I_{d_{2}}$;
  5. (v)
    trace-preserving if $\operatorname{\textsf{tr}}(\Phi(X))=\operatorname{\textsf{tr}}(X)$ for all $X\in\mathbb{M}_{d_{1}}$;
  6. (vi)
    unital quantum channel if $\Phi$ is a unital and trace-preserving CP map.
@@
Let $\langle X,Y\rangle:=\operatorname{\textsf{tr}}(X^{*}Y)$ be the Hilbert-Schmidt inner product on $\mathbb{M}_{d}$. Then the dual of a linear map $\Phi:\mathbb{M}_{d_{1}}\to\mathbb{M}_{d_{2}}$ is the unique linear map $\Phi^{*}:\mathbb{M}_{d_{2}}\to\mathbb{M}_{d_{1}}$ such that $\operatorname{\textsf{tr}}({\Phi(X)}^{*}Y)=\operatorname{\textsf{tr}}(X^{*}\Phi^{*}(Y))$ for all $X\in\mathbb{M}_{d_{1}}$ and $Y\in\mathbb{M}_{d_{2}}$. Denote the space of all linear maps from $\mathbb{M}_{d_{1}}$ to $\mathbb{M}_{d_{2}}$ by $\mathscr{B}({\mathbb{M}_{d_{1}},\mathbb{M}_{d_{2}}})$. In quantum information theory, the Choi-Jamiołkowski isomorphism [Cho75, Jam72], commonly known as channel-state duality, illustrates the connection between quantum channels and quantum states.
Theorem 2.2 (Choi-Jamiołkowski [Cho75, Jam72]).
@@
    $\Phi$ is Hermitian preserving if and only if $\mathscr{J}(\Phi)$ is Hermitian.
@@
Corollary 2.3 (Choi-Kraus [Cho75, Kra71]).
@@
The mixed unitary rank of $\Phi$ is the minimal $\ell$ for which such decomposition exists. It is immediate that, Choi-rank of $\Phi$ is less than or equal to the mixed unitary rank of $\Phi$. In [Bus06], it is proved that the mixed unitary rank of $\Phi$ is at most the square of the Choi-rank of $\Phi$, which is further less than or equal to $d^{4}$. For more details on mixed unitary rank, see [CKLi22].
It is well-known that every unital quantum channel on $\mathbb{M}_{2}$ is mixed unitary [Wat18, KuMa87]. This is not true in general for $d\geq 3$. For ready reference we have the following example.
@@
[Wat18, Example 4.3] (Holevo-Werner channel) Let $\Phi:\mathbb{M}_{3}\to\mathbb{M}_{3}$ be the unital quantum channel given by
$\Phi(X):=\frac{tr(X)I_{3}-X^{T}}{2},\quad\penalty 10000\ \penalty 10000\ \forall X\in\mathbb{M}_{3}.$
Then $\Phi$ is not a mixed unitary.
Let $\mathscr{H}(d)$ and $\mathscr{MU}(d)$ be the set of all Hermitian-preserving linear maps on $\mathbb{M}_{d}$ and the collection of all mixed unitary-channels on $\mathbb{M}_{d}$, respectively. The Choi-Jamiołkowski isomorphism $\mathscr{J}$ (c.f. 2.2) establishes a bijective correspondence between $\mathscr{B}({\mathbb{M}_{d},\mathbb{M}_{d}})$ and $\mathbb{M}_{d}\otimes\mathbb{M}_{d}$. Define an inner product on $\mathscr{B}({\mathbb{M}_{d}}):=\mathscr{B}({\mathbb{M}_{d},\mathbb{M}_{d}})$ by
@@
Then, $\mathscr{J}$ becomes an isometric isomorphism between two Hilbert spaces. Furthermore, we have the following:
@@
$(i)$ Since the set set of unitary channels is compact and $\mathscr{MU}(d)$ is its convex hull it is also compact. Clearly it is closed under compositions.
@@
Conversely, assume that $\langle\Gamma,\Phi\rangle\geq 0$ for all unital and trace-preserving $\Gamma\in\mathscr{MU}(d)^{\circ}$. Let $\delta_{d}(X)=\frac{1}{d}\operatorname{\textsf{tr}}(X)I_{d},\penalty 10000\ \penalty 10000\ \forall X\in\mathbb{M}_{d}$, and consider the sets,
$\mathcal{M}$ $=\{\Psi-\delta_{d}:\Psi\in\mathscr{MU}(d)\},$

### v2

@@ Abstract.
It is established that both discrete and continuous semigroups of unital quantum channels are eventually mixed unitary. This result is novel even for the subclass of Schur maps and stands in sharp contrast to the resolution of the asymptotic quantum Birkhoff conjecture by Haagerup and Musat, who demonstrated that tensor powers of some unital quantum channels maintain a persistent positive distance from the set of mixed unitary channels. Remarkably, our results show that this gap vanishes in finite time when considering ordinary powers within a semigroup.
Building on this, we define the mixed unitary index of a unital quantum channel as the minimum time (or power) beyond which all subsequent maps become mixed unitary. We demonstrate that for any fixed dimension $d\geq 3$, there is no universal upper bound for this index. Furthermore, we observe that if a continuous semigroup is not mixed unitary at some $t>0$, it remains non-mixed unitary for all $t$ sufficiently close to the origin. Finally, we investigate quantum dynamical semigroups where mixed unitarity is restricted to specific families, such as Weyl or diagonal unitaries. We show that Schur semigroups of correlation matrices eventually become mixtures of rank-one correlation matrices, and we characterize the generators of Schur semigroups that remain within this set for all $t\geq 0$.
@@
Completely positive maps, quantum channels, semigroup, mixed unitary channels, Schur channels
@@ 1. Introduction
For a natural number $d$, let $\mathbb{M}_{d}$ denote the algebra of all $d\times d$ complex matrices. In this article, we will be restricting ourselves to studying unital quantum channels (doubly stochastic maps) from $\mathbb{M}_{d}$ to $\mathbb{M}_{d}$. A unitary quantum channel on $\mathbb{M}_{d}$ is a quantum channel which is given by $X\mapsto U^{*}XU$, for all $X\in\mathbb{M}_{d}$ and for some unitary $U\in\mathbb{M}_{d}$. Unitary quantum channels are extreme in the convex set of unital quantum channels on $\mathbb{M}_{d}$. A quantum channel is said to be a mixed unitary if it is in the convex hull of the set of all unitary quantum channels, or equivalently (due to finite dimensionality) it is a convex combination of unitary channels. That is, a quantum channel $\Phi:\mathbb{M}_{d}\to\mathbb{M}_{d}$ is a mixed unitary if there exist $\{\lambda_{j}\}_{j=1}^{\ell}\subseteq[0,1]$ and unitaries $\{U_{j}\}_{j=1}^{\ell}\subseteq\mathbb{M}_{d}$ for some $l\in\mathbb{N}$ such that $\sum_{j=1}^{\ell}\lambda_{j}=1$ and
@@
Mixed unitary channels play a significant role in quantum information theory for several key reasons. They provide a rich class of examples that capture key properties of general quantum operations [34]. By virtue of their structural simplicity, they remain highly useful for theoretical analysis, often proving more tractable than arbitrary quantum channels. These channels emerge naturally in multiple contexts, including majorization frameworks based on Hermitian operators [1], as well as in quantum cryptographic protocols involving state encryption [2, 18, 3].
Taking inspiration from the classical Birkhoff-von Neumann theorem, which realizes all doubly stochastic maps as convex combinations of permutations, one may suspect that perhaps every unital quantum channel is mixed unitary. A well-known result [8, Proposition 2.6], [26, Theorem 1] establishes this for $d=2$. However, this property fails for $d\geq 3$ as first demonstrated by Tregub [36] and Kümmerer-Maassen [25]. See also [8, 26, 29, 39]. Additionally, determining whether a given unital channel is mixed unitary or not is known to be NP-hard [27].
@@
$\delta_{d}(X):=\frac{\mbox{trace}(X)}{d}I_{d},\penalty 10000\ \penalty 10000\ X\in\mathbb{M}_{d}.$ (1.2)
@@
A famous theorem of Watrous [38], says that there is a ball of positive radius around $\delta_{d}$ such that all unital channels in it are mixed unitary. As a consequence, any sequence of unital channels converging to the completely depolarizing channel eventually consists of only mixed unitary channels. This was taken note of in the context of powers of primitive unital channels in Section 4 of [17]. They define mixed unitary time (or index) of a unital quantum channel as a threshold $k$ such that all subsequent powers beyond which are mixed unitary. Moreover they show that such a threshold exists for primitive unital channels and obtain an upper bound for it.
@@
In a more recent work, Kribs et al., [23] extend this idea very significantly. Note that the completely depolarizing channel can be viewed as the unique trace-preserving conditional expectation map from $\mathbb{M}_{d}$ to the $C^{*}$-algebra $\mathbb{C}I_{d}$ of scalars. Fix a unital $C^{*}$-subalgebra $\mathcal{A}$ of $\mathbb{M}_{d}$ and let $\mathbf{E}_{\mathcal{A}}$ be the unique trace-preserving conditional expectation to it. It is a unital quantum channel. In [23] it is observed that unital quantum channels sufficiently close to $\mathbf{E}_{\mathcal{A}}$ and fixing elements of $\mathcal{A}$ are mixed unitary. Armed with this result, they show that if $\Phi$ is a unital quantum channel on $\mathbb{M}_{d}$, then there exists some $k\in\mathbb{N}$ such that $\Phi^{k}$ is mixed unitary [23, Theorem 22]. However, $\Phi^{k}$ fixes elements of an algebra $\mathcal{A}$, does not imply that $\Phi^{k+1}$ does the same. For this reason, they couldn’t show that $\{\Phi^{n}\}_{n\geq 0}$ has finite mixed unitary time. We overcome this difficulty in Theorem 3.9. In [16], Haagerup and Musat showed that in general tensor powers of unital quantum channels can maintain a positive distance from the set of mixed unitary channels. This settled the asymptotic quantum Birkhoff conjecture in the negative. In striking difference to this, here we see that ordinary powers of unital quantum channels become mixed unitary in finite time! This raises the natural question as to for fixed $d$, whether there exists a bound $N_{d}$ such that mixed unitary indices of all unital quantum channels on $\mathbb{M}_{d}$ are bounded by $N_{d}$. We answer this question in the negative in Theorem 4.15. We get it as a consequence of our analysis of continuous (one parameter) semigroups of unital quantum channels.
@@
A deep study of continuous semigroups of mixed unitary channels was carried out long time ago by Kümmerer and Maassen in [25]. Their goal was to understand Markov dilations of such semigroups. In the process they characterized generators of quantum dynamical semigroups which remain mixed unitary for all times. We obtain a different characterization of such generators (Theorem 5.2). We show that even in continuous time, semigroups of unital channels eventually become mixed unitary. In other words, they all have finite mixed unitary index (Theorem 4.12). We may note that there are no classical analogues for this kind of statements as classically all doubly stochastic matrices are convex combinations of permutation matrices.
As a special case of results above we see that if $c_{jk}]\in\mathbb{M}_{d}$ is a correlation matrix (positive semidefinite with diagonal entries equal to one), its Schur powers $\{[c_{jk}^{n}]:n\geq 0\}$ are eventually mixed rank-one matrices, that is, they are convex combinations of complex rank one correlation matrices and similar result holds for continuous Schur semigroups (Theorem 6.10). Furthermore, we provide a complete characterization of the generators $A=[a_{jk}]$, for which the correlation matrices $e^{ta_{jk}}]$ are mixed rank-one for all times $t\geq 0$ (See Theorem 6.11).
@@
We organize the paper as follows: In Section 2, we recall the necessary definitions and basic results required in the subsequent sections, and fix some notations. Section 3, investigates the mixed unitarity for a discrete semigroup $\{\Phi^{n}\}_{n\in\mathbb{N}}$ generated by an unital quantum channel $\Phi$ on $\mathbb{M}_{d}$. We first establish that every trace-preserving *-automorphism on a $C^{*}$-subalgebra $\mathcal{A}$ of $\mathbb{M}_{d}$ is necessarily of the form $\mathrm{Ad}_{U}$ for some $U\in\mathbb{M}_{d}$ (See Lemma 3.5). This helps us to prove that any unital quantum channel on $\mathbb{M}_{d}$ is asymptotically a conditional *-automorphism (Theorem 3.8). Building on this, Theorem 3.9 provides a key result of this work: “Any discrete semigroup $\{\Phi^{n}\}_{n\in\mathbb{N}}$ of unital channels is eventually mixed unitary.”
@@
Section 4 extends this analysis to continuous one-parameter semigroups $(\Phi_{t})_{t\geq 0}$ of unital quantum channels on $\mathbb{M}_{d}$. We establish an analogous result in Theorem 4.12, proving that these semigroups are also eventually mixed unitary. Furthermore, we demonstrate in Theorem 4.15 that there is no universal upper bound for the mixed unitary index in a fixed dimension $d\geq 3$. That is, no $N\geq 1$ exists such that $\Phi^{N}$ is a mixed unitary channel for every unital channel $\Phi$ on $\mathbb{M}_{d}$.
@@
In Section 5, we examine the local behavior of a continuous semigroup of a unital quantum channels $\{\Phi_{t}\}_{t\geq 0}$ near origin. First, a characterization theorem is established for such a semigroup on $\mathbb{M}_{d}$ to be mixed unitary for all times (Theorem 5.2). Using this Theorem, we prove that every continuous one parameter semigroup $\{\Phi_{t}\}_{t\geq 0}$ of unital quantum channels on $\mathbb{M}_{d}$ is either mixed unitary for all times or there is a time $t_{0}$ such that $\Phi_{t}$ is never mixed unitary for all times $t\in(0,t_{0})$ (Theorem 5.4). We end the Section with several examples of unital quantum channels with some explicit computations.
In the last section, we first discuss mixed Weyl-unitary channels on $\mathbb{M}_{d}$. We provide a complete structure theorem (Theorem 6.4) for a one-parameter semigroup $\{\Phi_{t}\}_{t\geq 0}$ of unital quantum channels on $\mathbb{M}_{d}$ to be Weyl covariant. We also introduce the notion of $G$-mixed unitary channel on $\mathbb{M}_{d}$ for any subgroup $G$ of $\mathbb{U}(d)$ and prove a characterization theorem (See Theorem 6.8) for a continuous one parameter semigroup of unital quantum channels on $\mathbb{M}_{d}$ to be $G$-mixed unitary for all times and close with our results on Schur maps.
@@ 2.1. Notations
Let $\mathbb{C}^{d}$ denote the $d$-dimensional complex Hilbert space with the inner product linear in the second variable and conjugate linear in the first. Let $\mathbb{M}_{d_{1}\times d_{2}}$ be the algebra of complex matrices of size $d_{1}\times d_{2}$, we use $\mathbb{M}_{d}$ to refer $\mathbb{M}_{d\times d}$. The collection of all positive (semidefinite) matrices and unitary matrices in $\mathbb{M}_{d}$ are denoted by $\mathbb{M}_{d}^{+}$ and $\mathbb{U}(d)$ respectively. For every $x\in\mathbb{C}^{d_{1}}$ and $y\in\mathbb{C}^{d_{2}}$, define the rank one linear map $|{x}\rangle\langle{y}|:\mathbb{C}^{d_{2}}\to\mathbb{C}^{d_{1}}$ by $|{x}\rangle\langle{y}|(z):=\langle y,z\rangle x$, for all $z\in\mathbb{C}^{d_{2}}$ Additionally, we let $\operatorname{\textsf{T}}=\operatorname{\textsf{T}}_{d}:\mathbb{M}_{d}\to\mathbb{M}_{d}$ be the transpose map (with respect to the standard orthonormal basis $\{e_{i}:1\leq i\leq d\}$).
@@
Let $\langle X,Y\rangle:=\operatorname{\textsf{tr}}(X^{*}Y)$ be the Hilbert-Schmidt inner product on $\mathbb{M}_{d}$, and $\left\lVert.\right\rVert_{2}$ be the corresponding Hilbert Schmidt norm on $\mathbb{M}_{d}$. Then the dual of a linear map $\Phi:\mathbb{M}_{d_{1}}\to\mathbb{M}_{d_{2}}$ is the unique linear map $\Phi^{*}:\mathbb{M}_{d_{2}}\to\mathbb{M}_{d_{1}}$ such that $\operatorname{\textsf{tr}}({\Phi(X)}^{*}Y)=\operatorname{\textsf{tr}}(X^{*}\Phi^{*}(Y))$ for all $X\in\mathbb{M}_{d_{1}}$ and $Y\in\mathbb{M}_{d_{2}}$. Denote the space of all linear maps from $\mathbb{M}_{d_{1}}$ to $\mathbb{M}_{d_{2}}$ by $\mathscr{B}({\mathbb{M}_{d_{1}},\mathbb{M}_{d_{2}}})$.
Trace-preserving completely positive maps are known as quantum channels. Any unital CP map $\Phi$ on $\mathbb{M}_{d}$, satisfies the Kadison-Schwarz inequality:
$\Phi(X^{*})\Phi(X)\leq\Phi(X^{*}X),\penalty 10000\ \forall X\in\mathbb{M}_{d}.$
From this, we see that every unital quantum channel on $\mathbb{M}_{d}$ is contractive with respect to the Hilbert-Schmidt norm. We make repeated use of this fact. In quantum information theory, the Choi-Jamiołkowski isomorphism [11, 20], commonly known as channel-state duality, illustrates the connection between quantum channels and quantum states.
Theorem 2.1 (Choi-Jamiołkowski [11, 20]).
@@
    $\Phi$ is Hermitian preserving (i.e., $\Phi(X^{*})=\Phi(X)^{*},\penalty 10000\ \forall X\in\mathbb{M}_{d}$) if and only if $\mathscr{J}(\Phi)$ is a Hermitian matrix.
@@
Corollary 2.2 (Choi-Kraus [11, 22]).
@@
The mixed unitary rank of $\Phi$ is the minimal $\ell$ for which such decomposition exists. It is immediate that, Choi-rank of $\Phi$ is less than or equal to the mixed unitary rank of $\Phi$. In [10], it is proved that the mixed unitary rank of $\Phi$ is at most the square of the Choi-rank of $\Phi$, which is less than or equal to $d^{4}$. For more details on mixed unitary rank, see [14].
It is well-known that every unital quantum channel on $\mathbb{M}_{2}$ is mixed unitary and this is not true in general for maps on $\mathbb{M}_{d}$ with $d\geq 3$. For ready reference we have the following example.
@@
[39, Example 4.3] (Holevo-Werner channel) Let $\Phi:\mathbb{M}_{3}\to\mathbb{M}_{3}$ be the unital quantum channel given by $\Phi(X):=\frac{1}{2}{(tr(X)I_{3}-X^{T})},\quad\forall X\in\mathbb{M}_{3}.$ Then $\Phi$ is not a mixed unitary.
Let $\mathscr{H}(d)$ and $\mathscr{MU}(d)$ be the set of all Hermitian preserving linear maps on $\mathbb{M}_{d}$ and the collection of all mixed unitary channels on $\mathbb{M}_{d}$, respectively. The Choi-Jamiołkowski isomorphism $\mathscr{J}$ (c.f. 2.1) establishes a bijective correspondence between $\mathscr{B}({\mathbb{M}_{d},\mathbb{M}_{d}})$ and $\mathbb{M}_{d}\otimes\mathbb{M}_{d}$. Define an inner product on $\mathscr{B}({\mathbb{M}_{d}}):=\mathscr{B}({\mathbb{M}_{d},\mathbb{M}_{d}})$ by
@@
Then, $\mathscr{J}$ becomes an isometric isomorphism between the Hilbert spaces $\mathscr{B}({\mathbb{M}_{d}})$ and $\mathbb{M}_{d}\otimes\mathbb{M}_{d}$. Furthermore, we have the following:
@@
$(i)$ Since the set of unitary channels is compact and $\mathscr{MU}(d)$ is its convex hull it is also compact. Clearly it is closed under compositions.
@@
Conversely, assume that $\langle\Gamma,\Phi\rangle\geq 0$ for all unital and trace-preserving $\Gamma\in\mathscr{MU}(d)^{\circ}$. Let $\delta_{d}$ be the depolarizing channel as in (1.2), and consider the sets,
$\mathcal{M}:$ $=\{\Psi-\delta_{d}:\Psi\in\mathscr{MU}(d)\},$